Details.

during the process of startup the bcu2k didn't come up a second try of operate was complete. The masterdiagnotic process didn't restarted and it can be seeing in the ocam not live. A restrart of the masterdiagnostic requires a power cycle of the bcu2k and set the bin again recovered the ocam

Test preset.

Fri Mar 18 19:55:18.563 2022 aos.PresetAO{FLAO}.complete - PresetAO{FLAO}[2956] - dx:nan dy:nan slNull: f1:0.0 f2:1.0 freq:391 mag:0 nBins:3 nModes:90 r0:nan snMode:0 strehl:0 ttMod:0-- JuanCarlosGuerra - 18 Mar 2022

ocam not live.

masterdiagnostic not up.

bcu2k power cycle.

reconfigure ocam 1x1 cropped.

Fri Mar 18 19:58:44.232 2022 aos.PresetAO{FLAO}.started - PresetAO{FLAO}[2957] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:16,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

star a bit defocused.

Close loop too much..Adsec failure.

Power off/on

Fri Mar 18 19:58:44.232 2022 aos.PresetAO{FLAO}.started - PresetAO{FLAO}[2957] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:16,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

record time to center the camera lens.?

how much time is converging...??

the camera lens didn;t move correctly.. what is it happening.

moving manually the camera lens and processes finish. CL posiiton:x=31.10,y=20.3 @341.

loop 200Hz. bin3.

loaded the disturbance

500Hz seeing 1.2"

Fri Mar 18 19:58:44.232 2022 aos.PresetAO{FLAO}.started - PresetAO{FLAO}[2957] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:16,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

center pupil it didn;t recover the diff in Y axis starts at -0.50 and it doesn't converge.

loop 390Hz

optimize+gopt.

3.39,1.49,0.10

gopt.

skipping frame ~10%

0.64,0.28,0.04

TN:20220318_205007.

stop loop... skipping frames.

pull out the star..

with the star out of the hot spot and illuminating one pupil....

set gain 0.1modi10

closed and skip frame after ok.

idl_129665 |INF| 793|2022-03-18 20:57:38.341710| MAIN > fsm_run_ao() start
idl_129665 |INF| 794|2022-03-18 20:57:39.392205| MAIN > fsm_run_ao() end
AdSecArbitrator_129747|INF| 667|2022-03-18 20:57:39.482141| COMMANDHANDLER > Command Run Ao (code 2014) successfully completed
AdSecArbitrator_129747|WAR| 668|2022-03-18 20:57:42.093732| MAIN > Skip frame detected!
20221803 20:57:58.9536 Alert: HighOrderOffload: HighOrderOffload vector ready

load a gain vector 2.8,0.6,0.10 ->loop pause->loadgain vector and the pupils are full illuminated...no skip frame. As expected.

load disturbance 200Hz, seeing 1.2"

converge in x axis but not in y.

optimize+gopt

skip frame HO1 at gain ~12

gopt on

0.56,0.26,0.03

no skip frames..

TN:_212631,_212717

diff x~0.3" not reducing...

gopt ko, goes from ko to ok...

ko

TN:_212926.

gopt ampl-> 20e-9

Fri Mar 18 21:32:34.744 2022 aos.PresetAO{FLAO}.started - PresetAO{FLAO}[2971] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:16.2,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

center pupil, the diffy is still stuck ~-0.4

a little push to y=10 and it converge. ....need to be investigated...

loop at 325Hz

optimize+gopt

failed first pass at ho1

manual gain: 4.10,1.10,0.05

gopt

no stable.. process failed three times.

Fri Mar 18 21:47:35.489 2022 aos.PresetAO{FLAO}.started - PresetAO{FLAO}[2975] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:16.2,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

gopt 30e-9

center pupil ok.

optimize+gopt.

4.13,2.29,0.10

_215449,215525

-215753

pause/resume/pause

saving data.

_215753 (8000 frames)

same seq

_215932 (8000 frames)

_220024

pause/resume/pause/resume

drop the signal to 1.9 pht/subap.

_220357

optimize+gopt

skip frame filaed

2nd try...

same failure.

apply gain manually:3.8,1.97,0.1

gopt ok....

0.69,0.38,0.06

_221145,_221207,

ncpa on

_221259,_221323

stop 3:15 pm

Topic revision: r2 - 19 Mar 2022, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback